Shift report 20011008 16:00-24:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Mon Oct 08 2001 - 00:28:16 EDT

  • Next message: Jens Jørgen Gaardhøje: "Re: run and analysis plan"

    ------------------------------------------------------------
    Date: Mon Oct  8 00:28:16 2001, Shift: 16:00-24:00
    Supervisor: R. Debbe, with: none
    Log book#9 pages 115 - 117
    
    Shift Summary:
    ------------------------------------------------------------
    Machine status: Now we run with beta* = 2 meters and 200 MHz
                    RF (rebucketed)
    
    Achievements: Two stores with with 480K trigger 1 MRS at 45
                  FFS at 12 BFS at 3
    
    Detector prob. Nothing new, T3 was turned off because it puts
                   too many mapping errors. Now that we run with
                   squeezed betas and rebucketed bunches the dead
                   time of the DAQ needs additional attention. It
                   may be time to use our nifty vertex trigger.
    
     
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Mon Oct 08 2001 - 00:28:41 EDT